find the general solutions for each of the following

  1. (3x2+9xy+5y2)dx-(6x2+4xy)dy =0 y(2)=-6
  2. x2dy/dx +xy=y3/x
1
Expert's answer
2021-07-06T15:42:27-0400

(1) Given equation is

"(3x^2+9xy+5y^2)dx-(6x^2+4xy)dy =0"


"(3x^2+9xy+5y^2)dx=(6x^2+4xy)dy"


"\frac{dy}{dx}=\frac{(3x^2+9xy+5y^2)}{(6x^2+4xy)}"


Putting y=vx, then "\frac{dy}{dx}=v+x\frac{dv}{dx}"


Then equation will be,

"v+x\frac{dv}{dx}=\frac{(3x^2+9vx^2+5v^2x^2)}{(6x^2+4vx^2)}"


"v+x\frac{dv}{dx}=\frac{(3+9v+5v^2)}{(6+4v)}"


"x\frac{dv}{dx}=\frac{(3+9v+5v^2)}{(6+4v)}-v"


"x\frac{dv}{dx}=\frac{(3+3v+v^2)}{(6+4v)}"


"\frac{(6+4v)dv}{(3+3v+v^2)}=\frac{dx}{x}"


Integrating both sides,


Integrating left hand side first,

"\int \frac{(6+4v)dv}{(3+3v+v^2)}"


Let "v^2+3v+3 = t"

"(2v+3)dv = dt"

Then integral will be,


"\int \frac{(6+4v)dv}{(3+3v+v^2)} = \int \frac{2 dt}{t} = 2ln|t| = 2ln|3+3v+v^2|"



Then, "2ln|3+3v+v^2| = ln|x| + ln|C|"


"(3+3v+v^2)^2 = Cx"

Putting v=y/x,


"(3+3(\frac{y}{x})+(\frac{y}{x})^2)^2 = Cx"


"(3x^2+3xy+y^2)^2 = Cx^5" is the solution of the equation.






(2) Given equation is, "x^2\frac{dy}{dx} +xy=\frac{y^3}{x}"


Then, "\frac{dy}{dx} +\frac{y}{x}=\frac{y^3}{x^3}"


Putting y=vx, we will get,


"v+x\frac{dy}{dx}+v = v^3"


"x\frac{dv}{dx} = v^3-2v"


"\frac{dv}{ v^3-2v} =\frac{dx}{x}"


Integrating both sides,

Integrating left-hand side first, "\int\frac{dv}{ v^3-2v}"


"\int\frac{dv}{ v^3-2v} = \int\frac{dv}{ v(v^2-2)}"

Doing partial fraction,

"\frac{1}{ v(v^2-2)} = \frac{A}{v}+\frac{Bv+C}{v^2-2}"

Solving For A, B and C, we get

"\frac{1}{ v(v^2-2)} = \frac{A(v^2-2)+Bv^2+Cv}{v(v^2-2)} = \frac{(A+B)v^2+Cv-2A}{v(v^2-2)}"


Comparing powers on both sides, we get

A+B=0, C=0 and -2A=1

"\implies A = -\frac{1}{2} , B = \frac{1}{2} , C=0"

Then,

"\frac{1}{ v(v^2-2)} = \frac{-1}{2v}+\frac{v}{2(v^2-2)}"

Now integrating it,


"\int\frac{1}{ v(v^2-2)} dv=\int \frac{-1}{2v}+\frac{v}{2(v^2-2)} dv"


"\int \frac{v}{2(v^2-2)} dv =\int \frac{2v}{4(v^2-2)} dv"

Taking "v^2-2 =t \implies 2vdv = dt"


"\int \frac{2v}{4(v^2-2)} dv = \int \frac{dt}{4t} = \frac{1}{4}ln|t|= \frac{1}{4}ln|v^2-2|"


So, integral of LHS will be,

"\int\frac{1}{ v(v^2-2)} dv=\int \frac{-1}{2v}+\frac{v}{2(v^2-2)} dv = -\frac{1}{2}ln|v|+\frac{1}{4}ln|v^2-2|"


So, we get,


"-\frac{1}{2}ln|v|+\frac{1}{4}ln|v^2-2| = ln|x|+ln|C|"


"-\frac{2}{4}ln|v|+\frac{1}{4}ln|v^2-2| = ln|x|+ln|C|"


"\frac{1}{4}ln|\frac{v^2-2}{v^2}| = ln|x| + ln|C|"


"\frac{v^2-2}{v^2} = Kx^4" where K is also some constant.


putting "v = \frac{y}{x}"


"\frac{y^2-2x^2}{y^2} = Kx^4" which is the required solution.